Exercise 10-20

Isabelle Abiassi operates a popular summer camp for elementary school children. Projections for the current year are as follows:
Sales revenue $7,860,000
Operating income $693,250
Average assets $3,952,000

The camp’s weighted-average cost of capital is 10%, and Isabelle requires that all new investments generate a return on investment of at least 14%. The camp’s current tax rate is 25%.

At last week’s advisory board meeting, Isabelle told the board that she had up to $50,000 to invest in new facilities at the camp and asked them to recommend some projects. Today the board’s president presented Isabelle with the following list of three potential investments to improve the camp facilities.
Playground Swimming Pool Gym
Incremental operating income $ 3,668 $ 5,096 $ 2,628
Average total assets 26,200 39,200 14,600
Calculate the return on investment, residual income, and economic value added for each of the three projects. (Enter negative amounts using either a negative sign preceding the number, e.g. -45 or parentheses, e.g. (45). Round Economic Value Added answer to 2 decimal places, e.g. 15.25 & all other answers to 0 decimal places, e.g. 15 or 15%.)
Playground Pool Gym
Return on Investment

%

%

%
Residual Income $

$

$

Economic Value Added $

$

$

LINK TO TEXT

LINK TO TEXT

LINK TO VIDEO

LINK TO VIDEO

Which of the three projects do you recommend Isabelle undertake?

Swimming PoolPlaygroundGymPlayground, Swimming Pool & GymPlayground & Swimming PoolPlayground & GymSwimming Pool & Gym
Answer #1

Ans:

playground

pool

gym

incremental operating income

2794

3384

2208

average assets

25400

37600

13800

1. Return on investment = incremental operating income / average assets * 100

playground

pool

gym

3668 / 26200 *100

= 14 %

5096 / 39200 *100

= 13 %

2628/ 14600 *100

= 18%

2. Residual income = incremental operating income - required rate of return * average assets

playground

pool

gym

3668 - .14* 26200

= $ 0

5096 - .14*39200

= ( $ 392)

2628 - .14*14600

=$ 584

3. Economic value added = incremental operating income net of tax - weighted average cost of capital * average assets

  

playground

pool

gym

3668(1-.25) - .10*26200

=2751- 2620

= $ 131

5096 (1-.25) -.10*39200

=3822 - 3920

=($ 98)

2628(1-.25) - .10 *14600

=1971 - 1460

= $ 511